### v8

Abstract
Due to the dynamic nature of the Semantic Web, version control is necessary to manage changes in widely used ontologies. Despite the long-standing recognition of ontology versioning (OV) as a crucial component of efficient ontology management, many approaches treat OV as similar to ontology matching (OM) and directly reuse OM systems for OV tasks. In this study, we systematically analyse similarities and differences between OM and OV and formalise the OM4OV pipeline to offer more advanced OV support. The pipeline is implemented and evaluated in the state-of-the-art OM system Agent-OM. The experimental results indicate that OM systems can be reused for OV tasks, but without necessary extensions, the current OM4OV pipeline can produce skewed measurements, poor performance in detecting update entities, and limited explainability of false mappings. To tackle these issues, we propose an optimisation method called the cross-reference (CR) mechanism, which builds on existing OM alignments to reduce the number of matching candidates and to improve overall OV performance.
keywords:
ontology matching, ontology versioning
category:
[…]
1 Introduction
Ontologies serve as the backbone of the Semantic Web, providing formal concept descriptions of shared concepts across various applications [14]. An ontology is not static, and the need for version control arises. While web data is dynamic, any ontology used must undergo periodic revisions to keep pace with the growth in domain knowledge, modifications to the application adaptation, or corrections to the shared conceptualisation [19]. For example, it is unrealistic to expect ontologies created in the 1990s to contain concepts such as “touchscreen”, “fingerprint sensor”, or “WiFi antenna” [15]. These modifications may cause undesirable deficiencies in artifacts that conform to or reuse the ontology being changed, leading to severe non-compliance and incompatibility issues in downstream tasks.
Ontology versioning (OV), also known as ontology evolution, aims to distinguish and recognise changes between different ontology versions. By doing so, data that conforms to the changed ontology, ontologies that reuse the changed ontology, or software that uses the changed ontology can apply the correct changes correspondingly [19]. Although OV plays an important role in ontology management, this area remains underexplored. Many studies [27, 26, 28, 7, 12, 34, 21, 22, 23, 30] treat OV as similar to ontology matching (OM) and assume OM systems can be reused for OV tasks. OM is a well-studied problem, so leveraging this deep research legacy for OV is attractive (we call this approach “OM4OV”).
In this paper, we assume the Web Ontology Language (OWL) [5] is used to present an ontology. We use “concept” and “entity” to refer to OWL classes or properties. We use “subsumption” to mean “subclass of” and “equivalence” to mean “equivalent” in OWL as formal concept inter-relations. We analyse the complementary relationship between the OM and OV tasks and provide a task formulation for the OM4OV pipeline. We implement the pipeline in a modern OM system to validate its effectiveness in unifying the OM and OV tasks. We investigate the links between these two tasks in practice and explore optimisations for OM4OV. To the best of our knowledge, this work is the first to systematically analyse the OM4OV pipeline and provide theoretical foundations and empirical evidence for its effectiveness and pitfalls. Specifically, our key contributions include:
  • We analyse the similarities and differences between OM and OV. In the OM4OV pipeline, we describe four disjoint subset alignments produced in the OV process: remain, update, add, and delete, each corresponding to a category of match or non-match. We provide formal definitions and formulations.
  • We implement the OM4OV pipeline in the state-of-the-art OM system Agent-OM [32] and its extension Agent-OV and evaluate the performance on the testbed derived from the Ontology Alignment Evaluation Initiative (OAEI) [8] datasets. We find that OM systems can be reused for OV tasks, but extensions are necessary to improve overall OV performance and its meaningful quality measurement.
  • We propose an optimised OM4OV pipeline. The new pipeline introduces a cross-reference (CR) mechanism, which builds on prior OM alignments to overcome the pitfalls of the naive OM4OV pipeline by reducing the number of matching candidates and filtering out ambiguous false mappings, thereby significantly improving overall OV performance.
The remainder of the paper is organised as follows. Section 2 reviews the current literature. Section 3 provides the task formulation for the OM4OV pipeline, while Section 4 evaluates the OM4OV pipeline and discusses several common pitfalls for improvement. We propose a cross-reference mechanism as a pipeline optimisation method and experimentally evaluate its performance in Section 5. Section 6 concludes the paper.
2 Related Work
[…]
Maintaining version information in the ontology can be labour-intensive and error-prone. Either extending the current schema or using change logs requires consistent updating over time. In most cases, this process is hand-crafted by the ontology engineer or requires human intervention (e.g. pre-defining a schema or creating a template for the change log). A manual process is more likely to make mistakes and fail to propagate changes to dependent artefacts. Also, in the real world, there is no guarantee that the ontology itself contains the complete version information or that it has a separate change log. In such cases, current approaches have limited capability to detect incorrect version information or to identify missing version information. These approaches rely on the version information contained in or attached to the ontology by manual methods. If such information is missing or incorrect, automated versioning of ontology concepts is necessary.
Reusing OM systems for OV tasks (OM4OV) is a lightweight and fully automatic approach. OM4OV is compatible with ontologies using different recording methods, including URIs or additional versioning triples, as well as ontologies that lack or do not include version information. The PromptDiff suite [27, 26, 28] is the canonical work that integrates different heuristic matchers and applies a matching-like comparison for OV. The concept of “matching” is later used in the H-Change methodology [7] to assess the assimilation of new concepts into the old ontology. While studies in [12, 34] highlight that changes in ontologies can lead to corresponding changes in ontology mapping sets, the authors in [21, 22, 23] state that updating matching alignments is the key to tracking ontology version changes. A recent work in [30] develops an ontology evolution maintenance framework that combines a set of algorithms to revalidate ontology alignments. However, existing naive OM4OV approaches directly use OM systems for OV and focus on developing new algorithms and architectures for ontology version control, and less attention has been paid to understanding the underlying differences between OM and OV. OM is not exactly the same as OV. OM determines matched entities between two different ontologies, while OV distinguishes add, delete, remain, and update entities over different versions of one ontology. The OM input consists of two distinct ontologies, whereas the OV input is expected to consist of two versions of a single ontology. The output of OM is a set of entity mappings, whereas the output of OV comprises four sets, each defined by the nature of the change made. These differences may affect the adaptability of reusing OM systems for OV tasks.
3 OM4OV Task Formulation
In this section, we draw concepts in OM to formalise OV tasks.
OM Preliminary: Given a source ontology ($O_{s}$) and a target ontology ($O_{t}$), the OM task is to find an alignment $A$ (i.e. a set of mappings) with respect to a given similarity threshold $s\in[0,1]$, defined as [10]: $A=\{(e1,e2,r,c)|e1\in O_{s},e2\in O_{t},s\leq c<1\}$, where $e1$ and $e2$ are ontology entities in $O_{s}$ and $O_{t}$ respectively, $r$ states the relation between $e1$ and $e2$ which can be equivalence ($\equiv$) or subsumption ($\subseteq$), and $c\in[0,1]$ is the level of confidence for the match $(e1,r,e2)$.
Definition 3.1.
Given an old version of an ontology ($O$) and a new version of the same ontology ($O^{\prime}$), an OV task can be considered as finding an alignment (A) that contains a collection of mappings over the two different versions. As illustrated in Figure 1, while an alignment for OM only considers matched entities, OV tracks both matched and non-matched entities. Furthermore, matched entities are composed of two subsets remain and update, while non-matched entities are composed of add and delete entities.
Figure 1: Overview of the OM4OV pipeline.
Definition 3.2.
An OV task can be formalised as finding two variants of an alignment $A_{match}$ and $A_{non-match}$ between ontology versions $O$ and $O^{\prime}$ with respect to a given similarity threshold $s\in[0,1]$. OV focuses only on the equivalence relation, and subsumption mappings are beyond the scope of versioning. We assume the relation $r$ is an equivalence and do not refer to the relation henceforward. Therefore, the OV task is defined as:
$A_{match}$ $=\{(e1,e2,c)|e1\in O,e2\in O^{\prime},s\leq c\leq 1\}$ (1)
$A_{non-match}$ $=\{(e1,e2,c)|e1\in O,e2\in O^{\prime},0\leq c<s\}$
Definition 3.3.
The remain and update entities are actually matched entities between different ontology versions. The remain entities can be considered to be the exact (trivial) equivalent string matches between different ontology versions, assuming that the naming convention does not change between different ontology versions. If the naming convention changes, we will classify ontology entities that have changed names according to the systematic convention as update entities, by modelling their confidence $c$ as a value greater than $s$ but less than 1. The update entities are, in general, the closest (non-trivial) near-equivalent matches between different ontology versions. Therefore, $A_{remain}(A\odot)$ and $A_{update}(A\otimes)$ are defined as:
$A_{match}=A\odot\cup A\otimes$ (2)
$A\odot=A_{match}(c=1)=\{(e1,e2,c)|e1\in O,e2\in O^{\prime},c=1\}$
$A\otimes=A_{match}(s\leq c<1)=\{(e1,e2,c)|e1\in O,e2\in O^{\prime},s\leq c<1\}$
Definition 3.4.
The add and delete entities are actually non-matched entities between different ontology versions. The add entities are the non-matched terms in $O^{\prime}$, which have no matches in $O$. Similarly, we can interpret the delete entities as non-matched entities in $O$. Without loss of generality, we consider the set of entities in the add and delete categories to be alignments by equivalently interpreting them. Therefore, $A_{add}(A\oplus)$ and $A_{delete}(A\ominus)$ are defined as:
$A_{non-match}=A\oplus\cup A\ominus$ (3)
$A\oplus=\{e2\in O^{\prime}\cap A_{non-match}\}=\{e2\in O^{\prime}\setminus A_{match}\}$
$A\ominus=\{e1\in O\phantom{{}^{\prime}}\cap A_{non-match}\}=\{e1\in O\phantom{{}^{\prime}}\setminus A_{match}\}$
4 OM4OV Evaluation
Now we turn to the OM4OV evaluation.
4.1 Dataset Construction
There is a dearth of benchmark datasets for evaluating OV. The Ontology Alignment Evaluation Initiative (OAEI) contains several datasets related to OM tasks, but none are specifically designed for OV tasks. We propose an approach to constructing synthetic OV datasets from OM datasets. Figure 2 illustrates the generation of OM4OV datasets.
Figure 2: Generation of OM4OV datasets.
{bracketenumerate}
The original OAEI datasets for OM provide two ontologies, the source $O_{s}$ and the target $O_{t}$. We choose one as the ontology for versioning and duplicate it into $O$ and $O^{\prime}$. Another ontology will be used later to provide references for update entities. We duplicate this ontology into $O_{r}$.
There are four possible entity changes in OV tasks: remain, update, add, and delete. We retrieve all ontology entities from $O$ or $O^{\prime}$ (since they are identical at this stage). Each ontology entity will be randomly assigned to one of these categories, and the four sets are pairwise disjoint by construction. In reality, there are limited entities that have replacements. To ensure that the percentage of update entities can reach 25% of the distribution, we prioritise the assignment of update entities and assign the other three categories until the maximum replacement or 25% ontology size is reached.
For each ontology entity $e$ assigned to a category, the corresponding dataset operations are described as follows. While there is no operation for remain entities, entities assigned to add and delete are required to delete related triples in $O$ or $O^{\prime}$. The update entities need to replace all related triples in $O^{\prime}$. We generate four corresponding versioning references based on the entity assignments. These files provide the ground truth for evaluation.
To ensure the replaced entity is reasonable and reflects real-world facts, we use equivalent classes or properties specified in reference.xml. For example, replace cmt:ProgramCommitteeChair with its equivalent class confof:Chair_PC. This operation is cascaded, meaning that we update not only the entity name but also its related triples, such as changing the semantic information from (cmt:ProgramCommitteeChair, rdfs:subClassOf, cmt:ProgramCommitteeMember) to (confof:Chair_PC, rdfs:subClassOf, confof:Person). While this procedure can be unlimited, we restrict it to extract only the triples directly related to the entity (i.e. a 1-hop subgraph) in the ontology $O_{r}$. For entities whose names are unique identifiers or codes (and not textually meaningful names), we retrieve their meaningful descriptions from annotation properties. For example, we extract an additional annotation triple (mouse:MA_0000270, rdfs:label, “eyelid tarsus”) for the entity mouse:MA_0000270. After extraction, we also need to update the URI from “confof” to “cmt” where the new entity URI will be cmt:Chair_PC and the replacement triple will be (cmt:Chair_PC, rdfs:subClassOf, cmt:Person). Finally, we add these new triples to the existing ontology $O^{\prime}$. Table 1 summarises the detailed steps of the dataset operation for update entities.
Table 1: Detailed steps of the dataset operation for update entities.
Step Dataset Operation
#1 Remove all the triples related to $e$ in $O^{\prime}$.
#2 Find the equivalent class/property $e_{r}$ in reference.xml.
#3 Extract all the triples related to $e_{r}$ in $O_{r}$.
(Optional) Extract annotation triples related to $e_{r}$ in $O_{r}$.
#4 Update the URI from $e_{r}$ to $e$ and create a new entity $e^{\prime}$.
#5 Add all the triples related to $e^{\prime}$ to $O^{\prime}$.
Unlike the original OAEI datasets used for OM, randomness ensures that the synthetic OAEI datasets for OV are different each time they are constructed. This suits the dynamic nature of OV, where the changes vary between different versions. For this reason, we consider the new OAEI datasets for OV more like a testbed, as they can simulate a variety of situations for OV tasks. For reproducibility, we use a fixed seed to control the randomness. Table 2 lists the detailed information of the selected OAEI track used for the OV testbed. The current version of the OV testbed contains a total of 11 distinct ontologies from three different OAEI tracks. The anatomy track contains two large ontologies, while the MSE track has two medium ontologies, and the conference track has 7 small ontologies. We exclude EMMO [9] in the MSE track because it uses a different naming convention than MaterialInformation [3] and MatOnto [18]. Figure 3 shows the number of entities in each track. The fixed seed for the OV testbed is set to 42.
Table 2: Selected OAEI tracks for the OV testbed.
Track Domain # of Ontologies Anatomy Human and Mouse Anatomy 2 Conference Research Conference 7 MSE Materials Science & Engineering 2
Figure 3: Number of entities.
4.2 System Implementation
We choose the matching systems Agent-OM and its extended version Agent-OV for the experiment. Agent-OM is one of the flagship OM systems powered by large language models (LLMs) and multi-agent architectures. Its foundation framework is designed for traditional OM tasks. We extend Agent-OM with the OM4OV pipeline so that it can be used for OV tasks. Inherited from Agent-OM, Agent-OV supports a wide range of LLMs, including commercial API-accessed LLMs OpenAI GPT [29] and Anthropic Claude [2], as well as open-source LLMs Meta Llama [24], Alibaba Qwen [1], Google Gemma [11], and ChatGLM [36]. Similarly to Agent-OM, similarity_thresholdsimilarity\_threshold and $top@k$ are the adjustable key hyperparameters for Agent-OV.
4.3 Evaluation Criteria
OM typically measures performance using precision, recall, and the F1 score. Given a gold standard reference (R) and a system-discovered alignment (A), Precision measures matching correctness and recall measures matching completeness, while F1 score offers a harmonic mean to balance correctness and completeness.
$Precision=\frac{|A\cap R|}{|A|}\quad Recall=\frac{|A\cap R|}{|R|}\quad F_{1}\ Score=\frac{2}{Precision^{-1}+Recall^{-1}}$ (4)
4.4 Experiment Results
Figure 4 shows the result of evaluating Agent-OM and Agent-OV on the OV testbed. We choose to use the Meta open-source model llama-3-8b for our experiments in this paper, eschewing the higher-end models which offer small improvements that may not justify the financial investment. We use embeddings from the same LLM model to further avoid expensive API calls and thereby provide an entirely free-to-use version of Agent-OV. The hyperparameter settings are similarity_threshold=0.90similarity\_threshold=0.90 and $top@k=3$ across all the system alignments generated. We set $top@k=3$ based on empirical experience that 3-5 works well for Agent-OM. We experiment with the alternative values of the similarity threshold later in this paper.
Figure 4: Evaluation of Agent-OM and Agent-OV on the OV testbed. Agent-OM only determines entity matches, while the extended Agent-OV determines four different categories of entity matches (remain or update) and non-matches (add or delete).
The results indicate that it is possible to unify the OM and OV tasks using the OM4OV pipeline. Both Agent-OM and Agent-OV can produce alignments with precision, recall, and F1 scores. However, we can see that Agent-OM consistently performs unusually well across all alignments. This is because the traditional OM system captures only matched entities between two ontologies, which in the OV task correspond to the remain and update entities. While remain entities commonly comprise the major proportion of entities over two different versions, the measure of remain entities dominates the overall evaluation and other categories are ignored. We believe this is a common situation that leads to misleading performance measurement when using the OM system in OV tasks. Agent-OV overcomes the skewed measure in Agent-OM. When decomposing the measure into four sub-measures, we can observe more informative matching performance across different OV tasks. In general, we observe that the matching performance is highest in remain, followed by add and delete, and relatively low in update. This trend is consistent across different tracks and ontologies. The measures for remain are typically very close to 100%. The measures for add and delete are generally good. Our system uses LLMs as the backend, and LLMs generally have strong background knowledge to detect non-matched entities. The measures for update show scope for improvement. This is because finding non-trivial alignments and appropriate similarity thresholds is not easy. Additionally, we observe a longer computation time in the Anatomy Track. Although Agent-OV has an optimisation module for the matching candidate selection process (inherited from Agent-OM), it is still insufficient for some OV tasks. There is a need to optimise the pipeline for OV in large-scale ontologies.
Figure 5 studies the effect of varying similarity thresholds in Agent-OV. We change the similarity threshold from 1.00 to 0.50 to evaluate its effect on matching performance. While OV entities are classified into four different categories, each category requires a corresponding sub-measurement. Similar to OM tasks, changes in hyperparameter settings in OV tasks lead to a trade-off between precision and recall. Moreover, the hyperparameter settings also influence the sub-measures. We find that the similarity threshold has no effect on detecting remain entities, but significantly affects the detection of update entities and slightly affects the detection of add and delete entities. Furthermore, detecting update entities and detecting add and delete entities are negatively correlated. Lower similarity thresholds can result in more update entities being detected, while higher similarity thresholds may find more add and delete entities.
Figure 5: Evaluation varying similarity thresholds in Agent-OV. We apply a 1-D Gaussian filter for smoothing. The standard deviation for a Gaussian kernel is 1 (i.e. $sigma=1$).
We further examine the false mappings produced by Agent-OV. Unlike OM tasks, we find that some of these false mappings are not actually untrue. Several examples are listed below.
{bracketenumerate}
False mappings in OV can result from flawed ontology design choices. In the following example, the reference shows that cmt:writtenBy is updated to cmt:isWrittenBy, but the OV system will consider this entity to remain unchanged because a false mapping (cmt:writtenBy, cmt:writtenBy, $\equiv$, 1) is detected in the remain subset. This issue is caused by inconsistent naming convention practice. In the CMT ontology, cmt:hasAuthor and cmt:writtenBy are different entities because one is targeting the paper (cmt:Paper, cmt:hasAuthor, cmt:Author) and one is targeting the review (cmt:review, cmt:writtenBy, cmt:reviewer). However, using cmt:writtenBy for the paper still makes sense (cmt:Paper, cmt:writtenBy, cmt:Author). Within one ontology, the meaning of two entities is too close to be distinguished and therefore leads to them being synonyms for each other. Ideally, we should avoid this type of ontology design. This example also demonstrates a unique benefit of using OM for OV, which could potentially assist in ontology design.
cmt:hasAuthor => cmt:writtenBy
cmt:writtenBy => cmt:isWrittenBy
False mappings in OV can create ambiguous “equivalent” relationships. In the following example, the reference shows that cmt:ConferenceChair can be updated to cmt:General_Chair to indicate a specific type of chair responsible for coordinating the conference. However, the OV system may predict that cmt:ConferenceChair is equivalent to cmt:Chair. This interpretation is not wrong, but follows a different ontology design pattern. It is vital to notice that the term “equivalent” in OV is weaker than that in OM. OV allows for roughly “equivalent” mappings. The entities mapped in OV can slightly alter their meanings in response to changes in domain understanding or evolution of natural language in the domain.
cmt:Chairman => cmt:Chair
cmt:ConferenceChair => cmt:General_Chair
False mappings in OV can arise from inappropriate settings of the similarity threshold. In the following example, if similarity_threshold(s)=0.95similarity\_threshold\ (s)=0.95, cmt:SubjectArea in $O$ will be assigned to delete entities as it does have matching entities; If similarity_threshold(s)=0.90similarity\_threshold\ (s)=0.90, cmt:SubjectArea in $O$ will be assigned to update entities as it has a matching entity cmt:Topic in $O^{\prime}$. Both results are valid because defining the boundary between matching and non-matching is context- and application-dependent. For example, the similarity threshold could be relatively higher in the biomedical domain to ensure that each term is unique, whereas the similarity threshold in the conference domain could be relatively lower to improve the interoperability of terminologies used in research conferences.
s = 0.95, cmt:SubjectArea => None
s = 0.90, cmt:SubjectArea => cmt:Topic
4.5 Discussion
The experimental results have shown that OM systems can be reused for OV tasks, but they hvae limitations and require necessary extensions.
{bracketenumerate}
Skewed Measurement: OM systems has skewed measurement of OV task because unchanged remain entities dominate in practice. OM measures are reusable, but these measures need to be extended into four sub-measures for add, delete, remain, and update performance. Within each sub-measure, the inherent precision-recall trade-off still holds. Across different sub-measures, they are not independent. This is because each change in a part of an alignment will cause other parts to change accordingly. For example, if a new mapping is found in $A\odot$ or $A\otimes$, then the number of mappings in $A\oplus$ and $A\ominus$ will be reduced by one each, and vice versa.
Update Pitfalls: OM systems are sensitive to changes in similarity thresholds and backend LLMs. For OV tasks, only update entities are sensitive to these changes, while remain, add, and delete are relatively stable across these hyperparameter settings. Poor update performance is common, and adjusting the hyperparameter settings does not help.
False Mappings: OM systems used for OV tasks can produce mappings that are false with respect to the ground truth reference, but that are true with respect to expert review. This can be caused by flawed ontology design choices, ambiguous “equivalent” relationships, or inappropriate settings of the similarity threshold. This is difficult for OM systems to recognise and generally requires human effort.
5 OM4OV Optimisation
Having established the current pitfalls of OM4OV, we explore optimisations for the pipeline. Often, ontology creators provide cross-references to other ontologies to enhance interoperability and facilitate integration. For example, a cross-reference between the CMT ontology and the Conference ontology is provided alongside the CMT ontology. Reusing the cross-references developed for OM tasks, we propose a novel mechanism to reduce the number of matching candidates and also to improve overall OV performance.
Figure 6 illustrates the cross-reference (CR) mechanism used in the OM4OV pipeline. We can see that, without using the cross-reference $O_{r}$, the matching candidates cover the range of $O\cup O^{\prime}$. This number can be significantly reduced by removing prior matches (i.e. $O\cap O_{r}\cap O^{\prime}$) and non-matches (i.e. $O\cap O_{r}\setminus O^{\prime}$ and $O^{\prime}\cap O_{r}\setminus O$). The prior matching will be incorporated into the final alignment, while the known non-matching regions will be removed entirely in the subsequent OV process. The OV process then only determines the posterior alignment.
Figure 6: Cross-reference mechanism.
In practice, the prior alignment usually contains a large number of remain entities and a small number of update entities. Matching performance is also enhanced by utilising these established mappings. Since prior matches are inferred from OM references validated by domain experts, they provide a solid ground truth for alignment within a specific domain. Removing the known non-matches from the OV process also simplifies the detection of the posterior alignment. The optimised OM4OV pipeline using the CR mechanism is defined as follows.
Definition 5.1.
Given a reference ontology ($O_{r}$) with an old version of an ontology ($O$) and a new version of the same ontology ($O^{\prime}$), two cross-references between $O$ and $O_{r}$ ($R_{or}$) and between $O^{\prime}$ and $O_{r}$ ($R_{o^{\prime}r}$) are defined as:
$R_{or}$ $=\{(e1,e3,c)|e1,e3\in O_{r}\cap O,s\leq c\leq 1\}$ (5)
$R_{o^{\prime}r}$ $=\{(e2,e3,c)|e2,e3\in O_{r}\cap O^{\prime},s\leq c\leq 1\}$
We can use $R_{or}$ and $R_{o^{\prime}r}$ to infer some known mappings between $O$ and $O^{\prime}$ before performing OV. We call these mappings a prior alignment ($A_{\pi}$). After subsequently performing OV, we have our posterior alignment ($A_{\pi^{*}}$). In this setting, $A_{match}$ in OV can be decomposed into two parts: $A_{match}=A_{\pi}\cup A_{\pi^{*}}$.
Definition 5.2.
$A_{\pi}$ can be directly inferred from the two cross-references $R_{or}$ and $R_{o^{\prime}r}$. The equivalence relation is transitive, so for $e1\in O$, $e2\in O^{\prime}$, and $e3\in O_{r}$, if $e1=e3$ and $e2=e3$ then $e1=e2$. $A_{\pi^{*}}$ aims to detect missing mappings from the cross-reference. None of these mappings would come from any pairwise intersection of $O$, $O_{r}$, and $O^{\prime}$ because $O\cap O_{r}\setminus O^{\prime}$ and $O^{\prime}\cap O_{r}\setminus O$ are pre-defined as non-matched entities, and the matched entities in $O\cap O_{r}\cap O^{\prime}$ have already been captured in the $A(\pi)$. As a result, $A_{\pi^{*}}$ can be defined within a smaller scope:
$A_{\pi}$ $=R_{or}\cap R_{o^{\prime}r}=\{(e1,e2,c)|e1,e2\in O\cap O_{r}\cap O^{\prime},s\leq c\leq 1\}$ (6)
$A_{\pi^{*}}$ $=\{(e1,e2,c)|e1\in O\setminus O_{r},e2\in O^{\prime}\setminus O_{r},s\leq c\leq 1\}$
Definition 5.3.
An ontology can have multiple cross-references available. In such cases, the prior reference becomes the union of all known cross-references ($R_{or1}$$R_{orn}$), and the ontology used in the posterior alignment ($O_{ra}$) becomes the union of all reference ontologies ($O_{r1}$$O_{rn}$). Therefore, $A_{\pi}$ and $A_{\pi^{*}}$ in multiple cross-references can be formulated as:
$A_{\pi}$ $=(R_{or1}\cap R_{o^{\prime}r1})\cup(R_{or2}\cap R_{o^{\prime}r2})\cup.\cup(R_{orn}\cap R_{o^{\prime}rn})$ (7)
$A_{\pi^{*}}$ $=\{(e1,e2,c)|e1\in O\setminus O_{ra},e2\in O^{\prime}\setminus O_{ra},c\geq s\}\ \text{where}\ O_{ra}=O_{r1}\cup O_{r2}\cup.O_{rn}$
Figure 7 compares the OV performance with and without the CR mechanism on the same OV testbed that we analysed in Section 4. We can see that the CR mechanism significantly enhances Agent-OV’s ability to detect update entities and slightly improves its performance in detecting add and delete entities, with no effect on remain entities. More specifically, performance improvement is mainly attributed to recall for update entities, whereas it is more evident in precision for add and delete entities.
Figure 7: Evaluation of the cross-reference (CR) mechanism on the OV testbed. Each radial chart shows measurements (precision, recall, and F1 score) for a change category over 12 datasets analysed around the perimeter, both with (orange) and without (blue) the CR mechanism applied.
Figure 8 compares the effects of different similarity thresholds in the CMT Ontology versioning. The results of the experiment show that the CR mechanism is less sensitive to the similarity threshold than the original OM4OV pipeline, suggesting that it can be helpful in scenarios where the optimal similarity threshold is unclear or difficult to determine.
Figure 8: Evaluation of the cross-reference (CR) mechanism on different similarity thresholds. We apply a 1-D Gaussian filter for smoothing. The standard deviation for a Gaussian kernel is 1 (i.e. $sigma=1$).
6 Conclusion
In this paper, we systematically analyse similarities and differences between OM and OV tasks and provide formal definitions and task formulations for OM4OV. The experimental results obtained from Agent-OM and Agent-OV demonstrate that the OM4OV pipeline is workable, though it has several limitations and requires extensions. To tackle these issues, we propose a novel optimisation method that uses a cross-reference mechanism to leverage reference alignments from OM for OV. This method (1) overcomes several pitfalls in using OM for OV tasks, (2) significantly reduces the number of matching candidates, and (3) improves overall OV performance.
While OM has a universal measure, our proposed measures for OV have four sub-measures. We plan to investigate a merged formula for OV sub-measures in the future. For example, using a harmonic mean to combine sub-measures. We also plan to apply the OM4OV pipeline in emerging domain ontologies that require consistent changes over time. For example, the smart building ontologies Brick Schema [4], RealEstateCore [16], and ASHRAE Standard 223P [25].
Declaration of GenAI and AI-Assisted Technologies
During the preparation of this work, no GenAI was used to produce the manuscript. However, GenAI was intrinsic to the research reported here and is properly identified throughout the work.
Appendix A Relevance to TGDK
This paper introduces a novel OM4OV pipeline to leverage ontology matching (OM) for ontology versioning (OV), along with a novel cross-reference mechanism for performance optimisation and a reusable benchmark testbed. OM is within the scope of the journal “Ontology & Entity Alignment”. Specifically, OV also falls within the scope of the Special Issue “Data Management for (Knowledge) Graphs” in “Integration and interoperability of graph data models”, “Integration of ML models (e.g. GNNs, LLMs) in graph data management systems”, and “Explainability, responsibility, and fairness in graph data management”.
Appendix B A Proposal for OAEI Ontology Versioning (OV) Track
B.1 Description
Participants attending other OAEI tracks are welcome to submit an entry to this track.
  • The OAEI OV track does not have its original datasets. The datasets come from other tracks and are being preprocessed using the OM4OV pipeline. Guest reviewers are invited to ensure that the generated datasets align with real-world facts.
  • The OAEI-OV track uses two different datasets in the preparation and official evaluation phases. The reference alignment is available to participants during the preparation phase but kept blind during the official evaluation phase.
  • The OAEI OV track has four sub-measures corresponding to remain, add, delete, and update entities. No universal measure is provided to ensure that the system’s performance is evaluated across different dimensions.
B.2 Modalities
Due to the diversity of platforms used, we accept direct submissions from participants.
B.3 Schedule
The schedule is available at the OAEI main page. The key dates for this track are:
  • Preparation & Execution Phase: test dataset (with reference alignments) released. Participants are encouraged to test their OM systems on these datasets and modify the system settings.
  • Evaluation Phase: official evaluation datasets (with no reference alignments) released. Participants are not encouraged to modify the system settings during this phase.
  • Before OM Workshop: preliminary results are released and open to public review.
  • After OM Workshop: final results are published.
B.4 Checklist
  • Register the system and provide the main contact email.
  • Submit system alignments corresponding to the blind datasets used in the evaluation phase.
  • Submit a model card to describe the hyperparameter settings of the system.