### v9

Abstract
Due to the dynamic nature of the Semantic Web, version control is necessary to manage changes in widely used ontologies. Despite the long-standing recognition of ontology versioning (OV) as a crucial component of efficient ontology management, many approaches treat OV as similar to ontology matching (OM) and directly reuse OM systems for OV tasks. In this study, we systematically analyse similarities and differences between OM and OV and formalise an OM4OV pipeline to offer more advanced OV support. The pipeline is implemented and evaluated in the state-of-the-art OM system Agent-OM. The experimental results indicate that OM systems can be effectively reused for OV tasks, but without necessary extensions, can produce skewed measurements, poor performance in detecting update entities, and limited explanation of false mappings. To tackle these issues, we propose an optimisation method called the cross-reference (CR) mechanism, which builds on existing OM alignments to reduce the number of matching candidates and to improve overall OV performance.
keywords:
ontology matching, ontology versioning
category:
[…]
1 Introduction
Ontologies serve as the backbone of the Semantic Web, providing formal concept descriptions of shared concepts across various applications [14]. An ontology is not static, and the need for version control arises. While web data is dynamic, any ontology in use must undergo periodic revisions to keep pace with the growth in domain knowledge, modifications to the application adaptation, or corrections to the shared conceptualisation [19]. For example, it is unrealistic to expect ontologies created in the 1990s to contain concepts such as “touchscreen”, “fingerprint sensor”, or “WiFi antenna” [15]. These modifications may cause undesirable deficiencies in artifacts that conform to or that reuse the ontology being changed, leading to severe non-compliance and incompatibility issues in downstream tasks.
Ontology versioning (OV), also known as ontology evolution, aims to distinguish and recognise changes between different ontology versions. By doing so, data that conforms to the ontology, ontologies that reuse the ontology, or software that uses the ontology can apply the correct changes correspondingly [19]. Although OV plays an important role in ontology management, this area remains underexplored. Many studies [26, 25, 27, 7, 12, 34, 21, 22, 23, 29] treat OV as similar to ontology matching (OM) and assume OM systems can be reused for OV tasks. OM is a well-studied problem, so leveraging this deep research legacy for OV is attractive. We call this approach “OM4OV”.
We analyse the complementary relationship between OM and OV tasks and provide a task formulation for the OM4OV pipeline. We implement the pipeline in a modern OM system to validate its effectiveness in unifying the OM and OV tasks. We investigate the links between these two tasks in practice and explore optimisations for OM4OV. To the best of our knowledge, this work is the first to systematically analyse the OM4OV pipeline and to provide theoretical foundations and empirical evidence for its effectiveness and pitfalls. Specifically, our key contributions include:
  • We analyse the similarities and differences between OM and OV. In the OM4OV pipeline, we describe four disjoint subset alignments produced in the OV process: remain, update, add, and delete, each corresponding to a subset of match or non-match in OM. We provide formal definitions and formulations.
  • We implement an OM4OV pipeline in the state-of-the-art OM system Agent-OM [32] and its extension Agent-OV and evaluate the performance on the testbed derived from the Ontology Alignment Evaluation Initiative (OAEI) [8] datasets. We find that OM systems can be reused for OV tasks, but extensions are necessary to improve both overall OV performance and its meaningful quality measurement.
  • We propose an optimised OM4OV pipeline. The new pipeline introduces a cross-reference (CR) mechanism, which builds on prior OM alignments to overcome the pitfalls of the naive OM4OV pipeline by reducing the number of matching candidates and filtering out ambiguous false mappings, thereby significantly improving overall OV performance.
In this paper, we assume the OWL 2 Web Ontology Language (OWL 2) [5] is used to present an ontology. We use both concept and entity to refer to OWL classes or properties. We use subsumption to mean rdfs:subClassOf or rdfs:subPropertyOf and equivalence to mean owl:equivalentClass or owl:equivalentProperty as formal class and property inter-relations.
The remainder of the paper is organised as follows. Section 2 reviews the current literature. Section 3 provides the task formulation for the OM4OV pipeline, while Section 4 evaluates the OM4OV pipeline and discusses several common pitfalls for improvement. We propose a cross-reference mechanism as a pipeline optimisation method and experimentally evaluate its performance in Section 5. Section 6 concludes the paper.
2 Related Work
Version control is recognised as a vital element in ontology management. Different versions of an ontology need to support interoperability so that changes do not impede the effective and sustainable use of the ontology.
[…]
Maintaining version information in the ontology can be labour-intensive and error-prone. Either extending the current standard 3-tuple schema or using change logs requires consistent updating over time. In most cases, this process is hand-crafted by the ontology engineer or requires human intervention (e.g. pre-defining a schema or creating a template for the change log). A manual process is more likely to make mistakes and fail to propagate changes to dependent artefacts. Also, in the real world, there is no guarantee that the ontology itself contains the complete version information or that it has a separate change log. In such cases, current approaches have limited capability to detect incorrect or missing version information. These approaches rely on the version information contained in or attached to the ontology by manual methods. If such information is missing or incorrect, automated versioning of ontology concepts is necessary.
Reusing OM systems for OV tasks (OM4OV) is a lightweight and fully automatic approach to detecting changes in ontologies. OM4OV is compatible with ontologies using different recording methods, including URIs or additional versioning triples, as well as ontologies that lack internal version information. The PromptDiff suite [26, 25, 27] is the canonical work that integrates different heuristic matchers and applies a matching-like comparison for OV. The concept of “matching” is later used in the H-Change methodology [7] to assess the assimilation of new concepts into the old ontology. Several studies [12, 34] note that updates to an ontology will generally create a requirement to update its existing alignments to external ontologies. Methods [21, 22, 23] are proposed to identify changes that affect alignments, and in some cases the framework is proposed to automate alignment updates [29], but these works do not address the discovery of alignments from an ontology to its own updated version.
While existing OM4OV approaches directly use OM systems for OV and focus on developing new algorithms and architectures for ontology version control, paying less attention to understanding the underlying differences between OM and OV. OM is not exactly the same as OV. OM determines matched entities between two different ontologies, while OV can distinguish add, delete, remain, and update changes over different versions of one ontology. OM input consists of two distinct ontologies, whereas OV input is expected to consist of two versions of a single ontology. The output of OM is a set of entity mappings, whereas the output of OV comprises four sets, each defined by the nature of the change made. These differences may affect the adaptability of reusing OM systems for OV tasks.
3 Task Formulation
Now we draw on concepts in OM to formalise OV tasks. Given an ontology $O$ we say that some entity $e\in O$ when $e$ is of interest to ontology matching or ontology versioning, typically representing a concept of the domain of interest being modelled by $O$. We consider an entity to be the full IRI of the reference in an ontology, or its equivalent abbreviation via a prefix, for example cmt:ProgramCommitteeChair. We consider an entity name to be the substring of an entity that provides a meaningful description of the entity (e.g. ProgramCommitteeChair), excluding any entity prefix or expanded URI (e.g. not cmt:ProgramCommitteeChair).
3.1 OM Preliminary
Given a source ontology ($O_{s}$) and a target ontology ($O_{t}$), the OM task is to find an alignment $A$ (i.e. a set of mappings) with respect to a given similarity threshold $s\in[0,1]$, defined as [10]: $A=\{(e_{1},e_{2},r,c)\mid e_{1}\in O_{s},e_{2}\in O_{t},s\leq c<1\}$, where $e_{1}$ and $e_{2}$ are ontology entities in $O_{s}$ and $O_{t}$ respectively, $r$ is the relation between $e_{1}$ and $e_{2}$ which is equivalence ($\equiv$) or subsumption ($\preceq$), and $c\in[0,1]$ is the level of confidence for the match $(e_{1},e_{2},r)$.
3.2 OM4OV Pipeline
Given an old version of an ontology ($O$) and a new version of the same ontology ($O^{\prime}$), the OV task can be considered as finding an alignment (A) from the old ontology to the new one. As illustrated in Figure 1, while an alignment for OM only mentions matched entities, OV needs to recognise both matched and non-matched entities. Furthermore, OV matched entities are comprised of two subsets: remain and update, while non-matched entities are comprised of add and delete entities. While some of these sets are comprised of entities, and others of mappings, when convenient henceforward, we will loosely refer to all of these as either alignments or sets of entities interchangeably.
Figure 1: Overview of the OM4OV pipeline. The sets are comprised of entities, with the matched intersection containing entities that are considered equivalent by an ontology alignment.
An OV task can be formalised as finding an alignment and a non-alignment, $A_{match}$ and $A_{non-match}$ respectively, between different ontology versions $O$ and $O^{\prime}$ with respect to a given similarity threshold $s\in[0,1]$. OV focuses only on the equivalence relation; subsumption mappings are beyond the scope of versioning.
Definition 3.1.
Given ontologies $O$ and $O^{\prime}$ and similarity threshold $s\in[0,1]$, the OV task is to find $A_{match}$ and $A_{non-match}$ as follows:
$A_{match}$ $=\{(e_{1},e_{2},c)\mid e_{1}\in O,e_{2}\in O^{\prime},s\leq c\leq 1\ \mbox{and there is no }(e_{1},e_{2},c^{\prime})\mbox{ with }c^{\prime}>c\}$ (1)
$A_{non-match}$ $=\{e\mid e\in O\cup O^{\prime}\mbox{and there is no }(e_{1},e_{2},c)\in A_{match}\mbox{ with }e\in\{e_{1},e_{2}\}\}$
For convenience, we will sometimes refer to the elements of $A_{match}$ as entities rather than alignments, by which we mean entities that are mentioned in mappings in $A_{match}$.
The remain and update entities are entities successfully matched over different ontology versions. The remain entities can be considered to be the exact (trivial) equivalent string matches between different ontology versions. On the other hand, if the convention used to represent an entity name changes (e.g. from EyelidTarsus to eyelid_tarsus), we would expect an OV system to classify the change as an update by modelling the confidence $c$ as a value greater than $s$ but less than 1. The update entities are, in general, the closest (non-trivial) near-equivalent matches between different ontology versions.
Definition 3.2.
$A_{remain}(A\odot)$ and $A_{update}(A\otimes)$ are defined as:
$A\odot=\{(e_{1},e_{2},c)\mid(e_{1},e_{2},c)\in A_{match}\mbox{ and }c=1\}$ (2)
$A\otimes=\{(e_{1},e_{2},c)\mid(e_{1},e_{2},c)\in A_{match}\mbox{ and }c<1\}$
We can see that $A_{match}=A\odot\cup A\otimes$
The add and delete entities are actually non-matched entities between different ontology versions. The add entities are the entities in $O^{\prime}$ which have no matches in $O$. Similarly, we can interpret the delete entities as entities in $O$ which have no matches in $O^{\prime}$.
Definition 3.3.
$A_{add}(A\oplus)$ and $A_{delete}(A\ominus)$ are defined as:
$A\oplus=\{e\mid e\in O^{\prime}\mbox{ and }e\in A_{non-match}\}$ (3)
We can see that $A_{non-match}=A\oplus\cup A\ominus$. It is expected that $A\oplus$ and $A\ominus$ are disjoint, but this would be violated when a faulty OM4OV pipeline determines that an equivalent pair of entities in each ontology does not match.
4 OM4OV Evaluation
Now we consider the question of how OM4OV can be evaluated.
4.1 Dataset Construction
There is a dearth of benchmark datasets for evaluating OV. The Ontology Alignment Evaluation Initiative (OAEI) contains several datasets related to OM tasks, but none are specifically designed for OV tasks. We propose an approach to constructing synthetic OV datasets from real-world OM datasets. Figure 2 illustrates the generation of OM4OV datasets.
Figure 2: Generation of OM4OV datasets from initial OM inputs of source ontology, target ontology, and reference alignment, and output versioning ontologies and reference alignments with four categories.
{bracketenumerate}
The original OAEI datasets for OM provide two ontologies, the source $O_{s}$ and the target $O_{t}$. We choose either one as the ontology for versioning and duplicate it into $O$ and $O^{\prime}$. Another ontology will be used later to provide references for synthetic update entities. We duplicate this ontology into $O_{r}$.
There are four possible entity changes in OV tasks: remain, update, add, and delete. We retrieve all ontology entities from $O$ or $O^{\prime}$ (since they are identical at this stage). Each ontology entity will be randomly assigned to one of these categories, and the four sets are pairwise disjoint by construction. In real-world OV, entity updates are relatively infrequent. We allow a user-defined parameter to fix the proportion of update entities in the synthetic dataset, defaulting to 25%. The update entities are randomly selected from the replacement corpus, so the number of update entities is limited by the size of the replacement corpus, which may be less than the user parameter value. For this reason, selection of update entities is prioritised over other categories, and excess entities are redistributed evenly over the other three categories.
For each ontology entity $e$ assigned to a category, the corresponding dataset operations are described as follows. While there is no operation for remain entities, entities assigned to add and delete are deleted from all triples that mention them in $O$ or $O^{\prime}$ respectively. Each update mapping requires that all occurrences of its domain entity in triples in $O^{\prime}$ are replaced by the corresponding co-domain entity of the mapping. We generate four corresponding versioning references based on the entity assignments. These files provide the ground truth for evaluation.
To ensure the replaced entity is reasonable and reflects real-world facts, we use equivalent classes or properties specified in reference.xml. For example, we could replace cmt:ProgramCommitteeChair with its equivalent class confof:Chair_PC. This operation is cascaded, meaning that we update not only the entity name but also its related triples, such as changing the semantic information from (cmt:ProgramCommitteeChair, rdfs:subClassOf, cmt:ProgramCommitteeMember) to (confof:Chair_PC, rdfs:subClassOf, confof:Person). While this procedure can be unlimited, we restrict it to extract only the triples directly related to the entity (i.e. a 1-hop subgraph) in the ontology $O_{r}$. For entities whose names are unique identifiers or codes (and not textually meaningful names), we retrieve their meaningful descriptions from annotation properties. For example, we extract an additional annotation triple (mouse:MA_0000270, rdfs:label, “eyelid tarsus”) for the entity mouse:MA_0000270. After extraction, we also need to update the URI from “confof” to “cmt” where the new entity URI will be cmt:Chair_PC and the replacement triple will be (cmt:Chair_PC, rdfs:subClassOf, cmt:Person). Finally, we add these new triples to the existing ontology $O^{\prime}$. Table 1 summarises the detailed steps of the dataset operation for update entities.
Table 1: Detailed steps of the dataset operation for update entities.
Step Dataset Operation
#1 Remove all the triples related to $e$ in $O^{\prime}$.
#2 Find the equivalent class/property $e_{r}$ in reference.xml.
#3 Extract all the triples related to $e_{r}$ in $O_{r}$.
(Optional) Extract annotation triples related to $e_{r}$ in $O_{r}$.
#4 Update the URI from $e_{r}$ to $e$ and create a new entity $e^{\prime}$.
#5 Add all the triples related to $e^{\prime}$ to $O^{\prime}$.
Unlike the original OAEI datasets used for OM, we introduce randomness to ensure that the synthetic OAEI datasets for OV are different each time they are constructed. This suits the dynamic nature of OV, where the changes vary between different versions. For this reason, we consider the new OAEI datasets for OV more like a testbed, as they can simulate a variety of situations for OV tasks. When reproducibility is required, we use a fixed seed to control the randomness. Table 2 details the OAEI tracks selected for the OV testbed. The current version of the OV testbed contains a total of 11 distinct ontologies from three different OAEI tracks. The anatomy track contains two large ontologies, while the MSE track has two medium ontologies, and the conference track has 7 small ontologies. We exclude EMMO [9] from the MSE track because it assigns opaque symbolic codes to entity names, using a naming convention that differs from MaterialInformation [3] and MatOnto [18], which use meaningful entity names. Figure 3 shows the number of entities in each ontology in each track.
Table 2: Selected OAEI tracks for the OV testbed.
Track Domain # of Ontologies Anatomy Human and Mouse Anatomy 2 Conference Research Conference 7 MSE Materials Science and Engineering 2
Figure 3: Number of entities.
4.2 System Demonstration
We choose the matching system Agent-OM and its extended version Agent-OV to demonstrate the OM4OV pipeline. Agent-OM is one of the flagship OM systems powered by large language models (LLMs) and multi-agent architectures. Its foundation framework is designed for traditional OM tasks. We extend Agent-OM with the OM4OV pipeline so that it can be used for OV tasks. Inherited from Agent-OM, Agent-OV supports a wide range of LLMs, including commercial API-accessed LLMs OpenAI GPT [28] and Anthropic Claude [2], as well as open-source LLMs Meta Llama [24], Alibaba Qwen [1], Google Gemma [11], and ChatGLM [36]. Similarly to Agent-OM, similarity_thresholdsimilarity\_threshold and $top@k$ are the adjustable key hyperparameters for Agent-OV.
4.3 Evaluation Criteria
OM typically measures performance using precision, recall, and the F1 score. Given a gold standard reference (R) and a system-discovered alignment (A), precision measures matching correctness and recall measures matching completeness, while F1 score offers a harmonic mean to balance correctness and completeness. OV can reuse these measures, but they need to be extended into four sub-measures for add, delete, remain, and update performance. While an OV method will trade off performance over each of these sets, an OV application problem may well prioritise accuracy on some over others.
$Precision=\frac{|A\cap R|}{|A|}\quad Recall=\frac{|A\cap R|}{|R|}\quad F_{1}\ Score=\frac{2}{Precision^{-1}+Recall^{-1}}$ (4)
4.4 Results
Figure 4 shows the performance of Agent-OM and Agent-OV on the OV testbed and sets a benchmark for further research progress. The fixed seed for the OV testbed is set to 42. We choose to use the Meta open-source model llama-3-8b for our experiments in this paper, eschewing the higher-end models which offer small improvements that may not justify the financial investment. We use embeddings from the same LLM model to further avoid expensive API calls and thereby provide an entirely free-to-use version of Agent-OV. The hyperparameter settings are similarity_threshold=0.90similarity\_threshold=0.90 and $top@k=3$ across all the system alignments generated. We set $top@k=3$ based on empirical experience that 3-5 works well for Agent-OM. We experiment with the alternative values of the similarity threshold later in this paper.
Figure 4: Evaluation of Agent-OM and Agent-OV on the OV testbed. Agent-OM only determines entity matches, while the extended Agent-OV determines four different categories of entity matches (remain or update) and non-matches (add or delete).
The results indicate that it is possible to unify the OM and OV tasks using the OM4OV pipeline. Both Agent-OM and Agent-OV can produce alignments with precision, recall, and F1 scores. However, we can see that Agent-OM consistently performs surprisingly well across all alignments. This is because the traditional OM system captures only matched entities between two ontologies, which in the OV task correspond to the remain and update entities. Because remain entities typically comprise the major proportion of entities over two different versions, the correctness of remain entities dominates the overall evaluation and other categories are de-emphasised. We believe this is a common situation that leads to misleading performance measurement when using an OM system in OV tasks.
Agent-OV overcomes the skewed measure in Agent-OM. By decomposing the measure into four sub-measures, we can observe more informative matching performance across different OV tasks. In general, we observe that the matching performance is highest in remain, followed by add and delete, and relatively low in update. This trend is consistent across different tracks and ontologies. The measures for remain are typically very close to 100%. The measures for add and delete are generally good. Our system uses LLMs as the backend, and LLMs generally have strong background knowledge to detect non-matched entities. The measures for update show scope for improvement. This is because finding non-trivial alignments and appropriate similarity thresholds is not easy. Additionally, we observe a longer computation time in the Anatomy Track. Although Agent-OV has an optimisation module for the matching candidate selection process (inherited from Agent-OM), it is still insufficient for some OV tasks. There is a need to optimise the pipeline for OV in large-scale ontologies.
Figure 5 studies the effect of varying similarity thresholds for the Mouse ontology in Agent-OV. We change the similarity threshold from 1.00 to 0.80 to evaluate its effect on matching performance. While OV entities are classified into four different categories, each category requires a corresponding sub-measurement. Similar to OM tasks, changes in hyperparameter settings in OV tasks lead to a trade-off between precision and recall. Moreover, the hyperparameter settings also influence the sub-measures. We find that the similarity threshold has no effect on detecting remain entities, but significantly affects the detection of update entities and slightly affects the detection of add and delete entities. Detecting update entities and detecting add and delete entities are negatively correlated, as is expected because a near-miss update alignment will fall back to a pair of non-match entities, one from each ontology. Therefore, lower similarity thresholds can result in more update entities being detected, while higher similarity thresholds may find more add and delete entities.
Figure 5: Evaluation varying similarity thresholds in Agent-OV. We apply a 1-D Gaussian filter for smoothing. The standard deviation for a Gaussian kernel is 1 (i.e. $sigma=1$).
We further examine the false mappings produced by Agent-OV. Unlike OM tasks, we find that some of these identified false mappings are not actually false. Several examples are listed below. {bracketenumerate}
False mappings in OV can result from flawed ontology design choices. In the following example, the reference shows that cmt:writtenBy is updated to cmt:isWrittenBy, but the OV system will consider this entity to remain unchanged because another mapping (cmt:hasAuthor, cmt:writtenBy, c) leads to a false mapping (cmt:writtenBy, cmt:writtenBy, 1) that is detected in the remain subset. In the CMT ontology, cmt:hasAuthor and cmt:writtenBy are different entities because one is targeting the paper (cmt:Paper, cmt:hasAuthor, cmt:Author) and one is targeting the review (cmt:review, cmt:writtenBy, cmt:reviewer). However, using cmt:writtenBy for the paper still makes sense (cmt:Paper, cmt:writtenBy, cmt:Author). Within one ontology, the meaning of two entities is too close to be distinguished and therefore leads to them being synonyms for each other. Ideally, we should avoid this type of ontology design. This example also demonstrates a unique benefit of using OM for OV, which could potentially assist in ontology design.
(cmt:writtenBy, cmt:isWrittenBy, c)
(cmt:hasAuthor, cmt:writtenBy, c)
False mappings in OV can create ambiguous “equivalent” relationships. In the following example, the reference shows that cmt:ConferenceChair can be updated to cmt:General_Chair to indicate a specific type of chair responsible for coordinating the conference. However, the OV system may predict that cmt:ConferenceChair is equivalent to cmt:Chair. This interpretation is not wrong, but follows a different ontology design pattern. It is vital to notice that the term “equivalent” in OV is weaker than that in OM. OV allows for roughly “equivalent” mappings. The entities mapped in OV can slightly alter their meanings in response to changes in domain understanding or evolution of natural language in the domain.
(cmt:ConferenceChair, cmt:General_Chair, c)
(cmt:Chairman, cmt:Chair, c)
False mappings in OV can arise from inappropriate settings of the similarity threshold. In the following example, if similarity_threshold=0.95similarity\_threshold=0.95, cmt:SubjectArea in $O$ will be assigned to delete entities as it does have matching entities; If similarity_threshold=0.90similarity\_threshold=0.90, cmt:SubjectArea in $O$ will be assigned to update entities as it has a matching entity cmt:Topic in $O^{\prime}$. Both results are valid because defining the boundary between matching and non-matching is context- and application-dependent. For example, the similarity threshold could be relatively higher in the biomedical domain to ensure that each term is unique, whereas the similarity threshold in the conference domain could be relatively lower to improve the interoperability of terminologies used in research conferences.
(cmt:SubjectArea, None, 0.95)
(cmt:SubjectArea, cmt:Topic, 0.90)
4.5 Discussion
The experimental results have shown that OM systems can be reused for OV tasks, but they have limitations and require necessary extensions. The following key points need to be taken into consideration when reusing OM systems for OV tasks. {bracketenumerate}
Skewed Measurement: OM systems exhibit skewed measurement of OV task because unchanged remain entities dominate in practice. OM measures are reusable, but these measures need to be extended into four sub-measures for add, delete, remain, and update performance. Within each sub-measure, the inherent precision-recall trade-off still holds. Across different sub-measures, they are not independent. This is because each change in a part of an alignment will cause other parts to change accordingly. For example, if a new mapping is found in $A\odot$ or $A\otimes$, then the number of mappings in $A\oplus$ and $A\ominus$ will be reduced by one each, and vice versa.
Update Pitfalls: OM systems are sensitive to changes in similarity thresholds and backend LLMs. For OV tasks, only update entities are sensitive to these changes, while remain, add, and delete are relatively stable across these hyperparameter settings. Poor update performance is common, and adjusting the hyperparameter settings does not help.
False Mappings: OM systems used for OV tasks can produce mappings that are false with respect to the ground truth reference, but that are true with respect to expert review. This can be caused by flawed ontology design choices, ambiguous “equivalent” relationships, or inappropriate settings of the similarity threshold. This is difficult for OM systems to recognise and generally requires human effort.
5 OM4OV Optimisation
Having established the current pitfalls of OM4OV, we explore optimisations for the pipeline. Often, ontology creators provide cross-references to alternative ontologies to enhance interoperability and facilitate integration. For example, a cross-reference between the Mouse ontology and the Human ontology is provided in the OAEI anatomy track. Reusing the cross-references developed for OM tasks, we propose a novel mechanism to reduce the number of matching candidates and also to improve overall OV performance.
Figure 6 illustrates the cross-reference (CR) mechanism used in the OM4OV pipeline. We can see that, without using the cross-reference $O_{r}$, the matching candidates cover the range of $O\cup O^{\prime}$. This number can be significantly reduced by removing prior matches (i.e. $O\cap O_{r}\cap O^{\prime}$) and non-matches (i.e. $O\cap O_{r}\setminus O^{\prime}$ and $O^{\prime}\cap O_{r}\setminus O$). The prior matching will be incorporated into the final alignment, while the known non-matching regions will be removed entirely in the subsequent OV process. The OV process then only determines the posterior alignment.
Figure 6: Illustration of the cross-reference mechanism.
In practice, the prior alignment usually contains a large number of remain entities and a small number of update entities. Matching performance is also enhanced by utilising these established mappings. Since prior matches are inferred from OM references validated by domain experts, they provide a solid ground truth for alignment within a specific domain. Removing the known non-matches from the OV process also simplifies the detection of the posterior alignment. The optimised OM4OV pipeline using the CR mechanism is defined as follows.
Definition 5.1.
Given a reference ontology ($O_{r}$), an old version of an ontology ($O$) and a new version of the same ontology ($O^{\prime}$), two cross-references between $O$ and $O_{r}$ ($R_{or}$) and between $O^{\prime}$ and $O_{r}$ ($R_{o^{\prime}r}$) are defined respectively as:
$R_{or}$ $=\{(e_{1},e_{3},c)\mid e_{1},e_{3}\in O_{r}\cap O,s\leq c\leq 1\mbox{ and there is no }(e_{1},e_{3},c^{\prime})\mbox{ with }c^{\prime}>c\}$ (5)
$R_{o^{\prime}r}$ $=\{(e_{2},e_{3},c)\mid e_{2},e_{3}\in O_{r}\cap O^{\prime},s\leq c\leq 1\mbox{ and there is no }(e_{2},e_{3},c^{\prime})\mbox{ with }c^{\prime}>c\}$
We can use $R_{or}$ and $R_{o^{\prime}r}$ to infer some known mappings between $O$ and $O^{\prime}$ before performing OV. We call these mappings a prior alignment ($A_{\pi}$). After subsequently performing OV, we have our posterior alignment ($A_{\pi^{*}}$). In this setting, $A_{match}$ in OV can be decomposed into two parts: $A_{match}=A_{\pi}\cup A_{\pi^{*}}$. $A_{\pi}$ can be directly inferred from the two cross-references $R_{or}$ and $R_{o^{\prime}r}$. By assuming the equivalence relation as extracted from the two cross-references is transitive, we have that $e_{1}\in O$, $e_{2}\in O^{\prime}$, and $e_{3}\in O_{r}$, if $e_{1}=e_{3}$ and $e_{2}=e_{3}$ then $e_{1}=e_{2}$. $A_{\pi^{*}}$ aims to detect missing mappings from the cross-reference. None of these mappings would come from any pairwise intersection of $O$, $O_{r}$, and $O^{\prime}$ because $O\cap O_{r}\setminus O^{\prime}$ and $O^{\prime}\cap O_{r}\setminus O$ are pre-defined as non-matched entities, and the matched entities in $O\cap O_{r}\cap O^{\prime}$ have already been captured in the $A(\pi)$. As a result, $A_{\pi^{*}}$ can be defined within a smaller scope.
Definition 5.2.
Prior alignment ($A_{\pi}$) and posterior alignment ($A_{\pi^{*}}$) are defined as:
$A_{\pi}$ $=R_{or}\cap R_{o^{\prime}r}=\{(e_{1},e_{2},c)\mid e_{1},e_{2}\in O\cap O_{r}\cap O^{\prime},s\leq c\leq 1\}$ (6)
$A_{\pi^{*}}$ $=\{(e_{1},e_{2},c)\mid e_{1}\in O\setminus O_{r},e_{2}\in O^{\prime}\setminus O_{r},s\leq c\leq 1\}$
An ontology can have multiple cross-references available. In such cases, the prior reference becomes the union of all known cross-references ($R_{or1}$$R_{orn}$), and the ontology used in the posterior alignment ($O_{ra}$) becomes the union of all reference ontologies ($O_{r1}$$O_{rn}$).
Definition 5.3.
$A_{\pi}$ and $A_{\pi^{*}}$ in multiple cross-references can be formulated as:
$A_{\pi}$ $=(R_{or1}\cap R_{o^{\prime}r1})\cup(R_{or2}\cap R_{o^{\prime}r2})\cup.\cup(R_{orn}\cap R_{o^{\prime}rn})$ (7)
$A_{\pi^{*}}$ $=\{(e_{1},e_{2},c)\mid e_{1}\in O\setminus O_{ra},e_{2}\in O^{\prime}\setminus O_{ra},c\geq s\}\ \text{where}\ O_{ra}=O_{r1}\cup O_{r2}\cup.O_{rn}$
Figure 7 compares the OV performance with and without the CR mechanism on the same OV testbed that we analysed in Section 4. We can see that the CR mechanism significantly enhances Agent-OV’s ability to detect update entities and slightly improves its performance in detecting add and delete entities, with no effect on remain entities. More specifically, performance improvement is mainly attributed to recall for update entities, whereas it is more evident in precision for add and delete entities.
Figure 7: Evaluation of the cross-reference (CR) mechanism on the OV testbed. Each radial chart shows measurements (precision, recall, and F1 score) for a change category over 12 datasets analysed around the perimeter, both with (orange) and without (blue) the CR mechanism applied.
Figure 8 compares the effects of different similarity thresholds for the Mouse ontology versioning. The results of the experiment show that the CR mechanism is less sensitive to the similarity threshold than the original OM4OV pipeline in Figure 5, suggesting that it can be helpful in scenarios where the optimal similarity threshold is unclear or difficult to determine.
Figure 8: Evaluation of the cross-reference (CR) mechanism on different similarity thresholds. The solid lines contribute to OV performance without the CR mechanism, while the dotted lines contribute to OV performance with the CR mechanism. We apply a 1-D Gaussian filter for smoothing. The standard deviation for a Gaussian kernel is 1 (i.e. $sigma=1$).
6 Conclusion
In this paper, we systematically analyse similarities and differences between OM and OV tasks and provide formal definitions and task formulations for OM4OV. The experimental results obtained from Agent-OM and Agent-OV demonstrate that the OM4OV pipeline is workable, though it presents several limitations. To tackle these issues, we propose a novel optimisation method that uses a cross-reference mechanism to leverage reference alignments from OM for OV. This method (1) overcomes several pitfalls in using OM for OV tasks, (2) significantly reduces the number of matching candidates, and (3) improves overall OV performance.
We observe three opportunities for future work. (1) While OM has a universal evaluation measure, our proposed measures for OV have four sub-measures. We plan to investigate a merged formula for OV sub-measures in the future, for example, using a harmonic mean to combine sub-measures. (2) In this paper, we limit the semantic analysis to noting changes in 1-hop subgraph relations that can be either ontology keywords (such as domain, range, and subclass) or domain concepts as object properties. A more complete analysis of structural evolution could employ description logic (DL)-based reasoning for subsumption checking, consistency validation, and detection of semantically equivalent expressions with only syntactic differences. This kind of analysis has not been previously explored in OM and is not addressed in this paper, but could be worthwhile further work. (3) We plan to apply the OM4OV pipeline in emerging domain ontologies that demonstrate regular changes over time. For example, it could be used to track version changes and improve interoperability in smart building ontologies, such as Brick Schema [4] and RealEstateCore [16] for which schema-interoperability is imperative [31].
Declaration of GenAI and AI-Assisted Technologies
During the preparation of this work, no GenAI was used to produce the manuscript. However, GenAI is intrinsic to the research reported here and is properly identified throughout the work.
Appendix A Relevance to TGDK
This paper introduces a novel OM4OV pipeline to leverage ontology matching (OM) for ontology versioning (OV), along with a novel cross-reference mechanism for performance optimisation and a reusable benchmark testbed. OM is within the scope of the journal “Ontology & Entity Alignment”. Specifically, OV also falls within the scope of the Special Issue “Data Management for (Knowledge) Graphs” in “Integration and interoperability of graph data models”, “Integration of ML models (e.g. GNNs, LLMs) in graph data management systems”, and “Explainability, responsibility, and fairness in graph data management”.
Appendix B A Proposal for OAEI Ontology Versioning (OV) Track
B.1 Description
Participants attending other OAEI tracks are welcome to submit an entry to this track.
  • The OAEI OV track does not have its original datasets. The datasets come from other tracks and are being preprocessed using the OM4OV pipeline. Guest reviewers are invited to ensure that the generated datasets align with real-world facts.
  • The OAEI-OV track uses two different datasets in the preparation and official evaluation phases. The reference alignment is available to participants during the preparation phase but kept blind during the official evaluation phase.
  • The OAEI OV track has four sub-measures corresponding to remain, add, delete, and update entities. No universal measure is provided to ensure that the system’s performance is evaluated across different dimensions.
B.2 Modalities
Due to the diversity of platforms used, we accept direct submissions from participants.
B.3 Schedule
The schedule is available at the OAEI main page. The key dates for this track are:
  • Preparation & Execution Phase: test dataset (with reference alignments) released. Participants are encouraged to test their OM systems on these datasets and modify the system settings.
  • Evaluation Phase: official evaluation datasets (with no reference alignments) released. Participants are not encouraged to modify the system settings during this phase.
  • Before OM Workshop: preliminary results are released and open to public review.
  • After OM Workshop: final results are published.
B.4 Checklist
  • Register the system and provide the main contact email.
  • Submit system alignments corresponding to the blind datasets used in the evaluation phase.
  • Submit a model card to describe the hyperparameter settings of the system.